# CMB-HD as a Probe of Dark Matter on Sub-Galactic Scales

### v2

@@ V Discussion
We also find that baryonic feedback effects change the shape of the matter power spectrum in ways that differ significantly from the change induced by non-CDM models. From the difference in the CMB-HD lensing signal alone, we can readily distinguish between them. This overcomes a significant challenge in efforts to measure non-CDM models; one often cannot tell if a given small-scale deviation from CDM is due to a non-CDM model or baryonic feedback effects.
@@ Acknowledgements.
The authors thank Boris Bolliet, Anthony Challinor, Francis-Yan Cyr-Racine, Rouven Essig, Gil Holder, Mathew Madhavacheril, David Marsh, Joel Meyers, Julian Munoz, Blake Sherwin, Anže Slosar, and Alexander van Engelen for useful discussions. The authors thank Miriam Rothermel for testing the code and Jupyter notebooks made public with this paper, and Keyi Chen for an early version of the bottom panel of Figure 6. AM and NS acknowledge support from DOE award number DE-SC0020441 and the Stony Brook OVPR Seed Grant Program. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC award HEP-ERCAPmp107.
Appendix A Comparison of Analytic Spectra Calculations to CAMB
@@
Appendix B Approximating the Lensing Signal-to-Noise Ratio per k Mode
@@
Table 6: We compare the CMB-HD lensing SNR obtained from Eq. 13 (columns labelled ”correct”) and from Eq. 42 (equivalent to Eq. 14 after summing over the redshifts and multipoles), where we apply the approximations described in Appendix B (columns labelled ”approx.”). We use either the full redshift range $z\in[0,1100]$ back to recombination, or the range $z\in[0,6]$ (used in Figures 2 and 7). We list the lensing SNRs for each individual spectrum $XY\in\{TT,TE,EE,BB,\kappa\kappa\}$ in the first five rows, and for the combination of all the spectra in the last row, calculated over the multipole ranges listed in Table 1. The SNRs obtained from Eq. 13 use the full covariance matrix (or the $XY\times XY$ block for individual spectra), including off-diagonal elements, with spectra obtained by integrating Eq. 8 over the given redshift range, as described in Section III.3. When using Eq. 42, we calculate the spectra and a corresponding diagonal covariance matrix within each redshift bin, and then take a sum over redshifts. Note that the first column contains the same information as Table 3. We emphasize that the approximate lensing SNR calculations are only used in Figures 2 and 7 for illustrative purposes.

### v3

@@ V Discussion
We also find that baryonic feedback effects change the shape of the matter power spectrum in ways that differ significantly from the change induced by non-CDM models. While a more comprehensive or complex feedback model may alter the details of the suppression, a broad range of numerical simulations that include hydrodynamics suggest a relatively universal behavior [54, 63, 64, 65, 66], which is not easily reproduced by alternate dark matter models. Since CMB-HD measures lensing accurately over a wide range of scales, it can readily detect this baryonic feedback effect at over $100\sigma$. Thus we expect in general CMB-HD to overcome a significant challenge in efforts to measure non-CDM models, which is that one often cannot tell if a given small-scale deviation from CDM is due to an alternative dark matter model or baryonic feedback effects. More refined and robust theory modeling of these effects would help to improve the accuracy of these forecasts and conclusions.
@@ Acknowledgements.
The authors thank Boris Bolliet, Anthony Challinor, Francis-Yan Cyr-Racine, Rouven Essig, Gil Holder, Mathew Madhavacheril, David Marsh, Joel Meyers, Julian Munoz, Blake Sherwin, Anže Slosar, and Alexander van Engelen for useful discussions. The authors thank Miriam Rothermel for testing the code and Jupyter notebooks made public with this paper, and Keyi Chen for an early version of the bottom panel of Figure 6. AM and NS acknowledge support from DOE award numbers DE-SC0020441 and DE-SC0025309 and the Stony Brook OVPR Seed Grant Program. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC award HEP-ERCAPmp107.
Appendix A Extragalactic Foregrounds
As discussed in Section III.1, our forecasts account for the residual extragalactic foreground signal that is expected to remain in the CMB-HD maps after foreground-subtraction procedures. Our extragalactic foreground model includes contributions from the thermal and kinetic Sunyaev-Zel’dovich effects (tSZ and kSZ, respectively), from the cosmic infrared background (CIB), and radio sources.
Appendix B Construction of the Covariance Matrix
As mentioned in Section III.1.4, we follow a similar approach to [38] to calculate the CMB-HD covariance matrix for the lensed or delensed CMB and the CMB lensing spectra, which we describe in detail below.
The blocks of the lensed and delensed covariance matrices are calculated analytically using the CLASS delens package [48] (unless otherwise specified). We denote each block by $\mathbb{C}_{\ell_{1}\ell_{2}}^{XY,WZ}$, where $XY$ and $WZ$ can each be one of $TT$, $TE$, $EE$, $BB$, and $\kappa\kappa$. This calculation accounts for Gaussian and non-Gaussian terms (i.e., diagonal and off-diagonal elements, respectively), including:
  • The Gaussian variance between different CMB spectra, given by
  • The Gaussian variance of the CMB lensing power spectrum, given by
    where $XY$ and $WZ$ are each one of $TT$, $TE$, $EE$.
    where
The total covariance for the CMB $\times$ CMB blocks is then the sum of Eq. 23 and either 25 or 26, $\mathbb{C}_{\ell_{1}\ell_{2}}^{XY,WZ}=\mathbb{C}_{\ell_{1}\ell_{2},\mathrm{G}}^{XY,WZ}+\mathbb{C}_{\ell_{1}\ell_{2},\mathrm{NG}}^{XY,WZ}$. We assemble the full five-by-five block covariance matrix by joining the CMB×$\times$×CMB blocks with the CMB×$\times$×lensing (Eq. 28) and lensing×$\times$×lensing blocks (Eq. 24 plus the simulation-based covariance matrix from [39]).
[…]
Appendix C Comparison of Analytic Spectra Calculations to CAMB
@@
Appendix D Approximating the Lensing Signal-to-Noise Ratio per k Mode
@@
Table 6: We compare the CMB-HD lensing SNR obtained from Eq. 7 (columns labelled ”correct”) and from Eq. 42 (equivalent to Eq. 8 after summing over the redshifts and multipoles), where we apply the approximations described in Appendix D (columns labelled ”approx.”). We use either the full redshift range $z\in[0,1100]$ back to recombination, or the range $z\in[0,6]$ (used in Figures 2 and 7). We list the lensing SNRs for each individual spectrum $XY\in\{TT,TE,EE,BB,\kappa\kappa\}$ in the first five rows, and for the combination of all the spectra in the last row, calculated over the multipole ranges listed in Table 1. The SNRs obtained from Eq. 7 use the full covariance matrix (or the $XY\times XY$ block for individual spectra), including off-diagonal elements, with spectra obtained by integrating Eq. 2 over the given redshift range, as described in Section III.3. When using Eq. 42, we calculate the spectra and a corresponding diagonal covariance matrix within each redshift bin, and then take a sum over redshifts. Note that the first column contains the same information as Table 3. We emphasize that the approximate lensing SNR calculations are only used in Figures 2 and 7 for illustrative purposes.